Circular No. 6693

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1996 F2 (SOHO) C. St. Cyr, Naval Research Laboratory, reports, for the SOHO-LASCO

Consortium (see IAUC 6685), the discovery by D. Biesecker and B. McCarty of another Kreutz sungrazer in SOHO-LASCO C3 coronagraph archival data. Measurements by St. Cyr have been reduced by G. V. Williams and are given on MPEC 1997-N05 in detail, the approximate discovery position being as follows:

1996 UT  α_{2000}  δ_{2000}
Mar. 23.981  0^{h}44.^{m}1  −0^{o}19′

This object was perhaps of m_{1} = 3–4, and it was also visible with a tail in a single C2 image on Mar. 25.42 UT at 5.2 R_{⊙} from the center of the sun. The orbit computed by the undersigned, also given in detail on MPEC 1997-N05, indicates perihelion passage on Mar. 25.66.

MARS M. Gaskell, Department of Physics and Astronomy, University of Ne-

braska, reports that possible dust activity was detected in the Ascraeus-Ceraunius region near the Martian evening limb on June 26.09 UT. The detection was made using a 0.20-m reflector at 475× and a red W25 filter. Similar observations on June 28.1, 29.1, 30.1 and July 1.1 suggested that the dust continued, roughly along the equator, to the Chryse region. R. McKim, British Astronomical Association, remarks that similar dust storms in the Ceraunius region were seen at the identical season in 1935 and 1978.

BL LACERTAE J. C. Noble and M. T. Carini, Western Kentucky University; H. R.

Miller, Georgia State University; and T. J. Balonek, K. Whitman and S. M. Davis, Foggy Bottom Observatory (FBO), Colgate Observatory, report: “Observations with the 1.1-m Hall (H) telescope at the Lowell Observatory and the 0.4-m telescope at FBO show BL Lac to be in a major outburst, with R magnitudes as follows: May 18, 14.04 ± 0.01 (FBO); 19, 13.95 ± 0.06 (H); 27, 14.09 ± 0.05 (H); 28, 14.17 ± 0.05 (H); 29, 13.80 ± 0.08 (H); June 4, 14.30 ± 0.01 (FBO); 11, 14.22 ± 0.02 (FBO); 15, 13.71 ± 0.02 (FBO); 16, 13.93 ± 0.02 (FBO); 23, 12.50 to 12.75 (FBO) over a 4-hr interval. Further observations are encouraged at all wavelengths.”

1997 July 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Brian G. Marsden